Hi how do you solve these two question with systems of linear equations?27. [R] Is the vector ? w the water (0) -- 1.0 -----0 4 28. [R] Is the vector ? 4 3
Answer #1

let 2 3 - 3 For some C,CEIR Then - 3 locitc2 34 +02 40, +302 -9, + C2 = 1 - 0 30, +C₂ = 3 44 +362 = 2.- O and By solving -a,and The solution By sewing which is impossible 23 are not equal There 9c2EIR. doanot sainty So span of 3 doesnot belongs to tsubstitute C2 = 2 in Eq -9-262) = 1 -C, +4 = 1 -=-3 C = 45 and solving q- By - 202 - 1 fit c2 1 -C2=2 TC2= -2 Substitute C₂ =1 and By Equation Solving 39+40=1 40+4C2 = 4 €97 -ci = -3 14-3 Substitute a=3 in eq - weget c2 = -2 By solving equatico O and